User Tools

Site Tools


operations:agenda2018_04_05

Chair: Patrick

Attended: Jonny, Atifur, Mas, Simin, Katie, Tiege, Jamie

Apologies: Arwin, Ross, Lucas

Absent:

Handover Issues

  • CRF106:
    • yg: Assigned module (HOB+0107) was full, but used HOB+0097 instead. Module allocation spreadsheet wasn't updated in time.
    • yg: Eremotectl connection dropped out, caused experiment to start late. Network reliability issue, Brett might take a look at it.
    • If eremotectl dies, can try restarting ercd daemon on pcfs machine.
    • If network connection completely drops out call Yg, and they can reboot router from their end.
  • R1836:
    • yg: mark5 time 0.2 second off from fs time and computer time, but delays were stable. Experiment already started, didn't want to introduce a clock jump so just left it. Probably due to ntp time small errors, should correct itself.
  • R4834:
    • ke: Unlabelled module (StreamStor1) and too dark to verify. Shouldn't be an issue now that Ke interior light fixed

General Issues

  • Ke interior light fixed. Can turn on at pdu1ke.phys.utas.edu.au. If stops working, probably bulb burned out, let Mick know (auscopeke mailing address).
  • Hobart26m drives major failure. Encoder has failed, in repair. Need to do new pointing once it's fixed. Out of action for at least this week.
  • slogit script is fixed on all machines
  • Use comments.py or comments.sh? comments.py is best if it works, but shell script is a backup
/home/www/auscope/opswiki/data/pages/operations/agenda2018_04_05.txt · Last modified: 2018/04/05 03:12 by Patrick Yates